OpenStudy (anonymous):

Cashews cost $2.50 per pound. Peanuts cost $5 per pound. Steven wants a mixture of cashews and peanuts that costs $4.50 per pound and weighs 10 pounds. How many pounds of cashews and how many pounds of peanuts are needed to make this mixture?

OpenStudy (anonymous):

So let's form some equations Let C = Cashews and P = Peanuts The weight of C and P is equal to 10 pounds. C + P = 10 The mixture costs $4.50 per pound which is $45 for 10 pounds. 2.5C + 5P = 45 Now you see that they are simultaneous equations. Use either elimination or substitution to solve them. I'm gonna use substitution 1. C + P = 10 2. 2.5C + 5P = 45 From equation 1, C = 10 - P Substitute into equation 2 2.5(10-P) + 5P = 45 25 - 2.5P + 5P = 45 2.5P=20 P=20/2.5 = 8 Now resubstitute into 1 to find C C + 8 = 10 C = 2

OpenStudy (anonymous):

Thank you for your help.